OpenStudy (anonymous):

Find the solution to the following system of equations: x+2y=6 and 2x+y=9.

OpenStudy (anonymous):

Step 1: Rearrange x + 2y = 6 x = 6 - 2y Step 2: Substitute 2 (6 - 2y) + y = 9 Step 3: Solve 12 - 4y + y = 9 12 - 3y = 9 3y = 12 - 9 3y = 3 y = 1 Step 4: Find x x + 2y = 6 x + 2 (1) = 6 x + 2 = 6 x = 4